A dentist's drill starts from rest. After 2.90 s of constant angular acceleration it turns at a rate of 2.75 ✕ 104 rev/min. (a) Find the drill's angular acceleration. (b) Determine the angle (in radians) through which the drill rotates during this period. rad
Answer #1

(a)
first convert rev/min to rad/s

ω = 2.75*104 rev/min = (2.75*104 rev/min)(2π rad/rev)(min / 60 s) = 2879.79 rad/s

α = (ω - ωo)/t

α = 2879.79 rad/s / (2.90 s)
α = 993.03 rad/s^2

(b)
θ = αt^2 / 2
θ = (993.03 rad/s^2)(2.90 s)2 / 2
θ = 4.175*103 rad